OpenStudy (anonymous):

the area is 21 what is width and length if length of a rectangle is 8centimers less than 5 times its width.

OpenStudy (anonymous):

W = width L = length L = 5w-8 Area = (5w-8)(w) 21= (5w-8)(w) 21= 5w^2 -8w 0 = 5w^2 -8w -21 = (5w + 7)(w- 3) w = -7/5 Or w= 3 But it can't be negative cause it is a width, so it can only be 3 Sub in 5w- 8 = 5(3) - 8 = 15 - 8 = 7 Check (5(3) - 8) (3) (Looking for area) = (15-3) (3) = (7)(3) 21

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

You are welcome :)

OpenStudy (anonymous):

wait what would be the length

OpenStudy (anonymous):

@hikat

OpenStudy (anonymous):

The length will be the sub in bit, because we know that the length = 5w - 8 So, we just need to sub in the width(which is 3) in, and the answer will be the length